# Emergent Response Planning in LLM

### v1

@@ Abstract
In this work, we argue that large language models (LLMs), though trained to predict only the next token, exhibit emergent planning behaviors: their hidden representations encode future outputs beyond the next token. Through simple probing, we demonstrate that LLM prompt representations encode global attributes of their entire responses, including structural attributes (response length, reasoning steps), content attributes (character choices in storywriting, multiple-choice answers at the end of response), and behavioral attributes (answer confidence, factual consistency). In addition to identifying response planning, we explore how it scales with model size across tasks and how it evolves during generation. The findings that LLMs plan ahead for the future in their hidden representations suggests potential applications for improving transparency and generation control.
@@ 1 Introduction
In this work, we argue that LLMs, though trained to predict only the next token, display emergent planning behaviors: their hidden representations encode their future outputs beyond just the next token. Specifically, we observe that LLM prompt representations encode interesting global attributes of their upcoming responses. We call this phenomenon response planning and classify these global attributes into three categories: structural attributes (response length, reasoning steps), content attributes (character choices in storywriting, multiple-choice answers at the end of response), and behavioral attributes (answer confidence, factual consistency).
We empirically identify response planning by training simple probes on LLM prompt representations to predict the global attributes of their upcoming responses. We find that these probes achieve non-trivial prediction accuracy, providing strong evidence that LLMs plan at least part of their entire response in advance as soon as they read the prompt. Through further ablation experiments, we find that planning abilities positively scale with fine-tuned model size, peak at the beginning and end of responses, share certain planning patterns across models, and exceed models’ self-reported awareness.
The contribution of our work is two-fold: (1) To our best knowledge, we introduce the first formal definition and framework of emergent response planning in LLM. (2) We demonstrate empirically that LLMs perform emergent response planning through systematic probing experiments across various attributes types and tasks, and investigate their properties. These findings shed light on LLMs’ internal mechanisms and suggest novel approaches for predicting and controlling outputs pre-generation, potentially enhancing model controllability.
@@ 2 Related Work
Prior works exploring response planning in LLM. Previous study have examined whether LLMs can anticipate beyond the next token. Future Lens (Pal et al., 2023) models token distributions beyond the immediate next token using linear approximation.  (Geva et al., 2023) studies how LLMs retrieve factual associations during generation, while  (Men et al., 2024) extends this to Blocksworld planning, suggesting LLMs consider multiple planning steps simultaneously.  (Pochinkov et al., 2024) finds that tokens at context-shifting positions may encode information about the next paragraph.  (Wu et al., 2024) hypothesizes LLMs’ lookahead capability and tests two mechanisms—pre-caching and breadcrumbs—in a myopic training setting.
While prior works examine relatively narrow aspects like predictions several tokens ahead or knowledge retrieval in specialized scenarios, our work delves deeper to reveal the broader response planning landscape of LLMs. We provide the first formal definition of response planning in LLM, investigate comprehensive planning attributes, and demonstrate planning capabilities across diverse real-world tasks.
@@ 3.2 Probing Setup
To study response planning in LLMs, we first design tasks $T=(p(\mathbf{x}),g(\mathbf{y}))$, consisting of a prompt distribution $p(\mathbf{x})$ eliciting key response attributes of interests $g(\mathbf{y})$ as probing targets. Next, we introduce the data collection pipeline for training probes.
Task design. The studied response attributes must be global, meaning they cannot be determined from the first response token and should ideally be distributed across the entire response. We focus on six tasks that elicit response attributes across three categories: structural, content, and behavioral.
@@
d Multiple-choice answers prediction.
@@
a Response length prediction.
c Character choices prediction.
e Factual consistency prediction.
b Reasoning steps prediction.
d Multiple-choice answers.
f Answer confidence prediction.
g Example fitting results for response length prediction.
h Example fitting results for reasoning steps prediction.
@@ 3.3 Experimental Details
We experiment with both instruction-tuned models (Llama-2-7B-chat, Llama-3-8B-Instruct, Mistral-7B-Instruct, and Qwen2-7B-Instruct) and their corresponding base models (Llama-2-7B, Llama-3-8B, Mistral-7B, and Qwen2-7B). See Appendix A.1 for model specifications and links.
@@ 4 Experimental Results
Insight 1: Models present emergent planning on structural, content, and behavioral attributes, which can be probed with high accuracy (Fig. 1).
Our in-dataset probing experiments (where probes are trained and tested on different splits of the same prompt dataset) reveal that both base and fine-tuned models encode structural, content, and behavioral attributes, with fine-tuned models showing superior performance. For structural attributes (response length and reasoning steps; Fig. 1a,  1b), fine-tuned models exhibit strong linear correlations with ground truth, clustering around $y=x$ (with example fitting results shown in Fig. 1g,  1h), while base models show weaker but positive correlations. For content and behavior attributes (character choices, multiple-choice answers, answer confidence, and factual consistency; Fig. 1c,  1d,  1e,  1f), both model types demonstrate robust classification performance above random baselines. These findings also suggest that models develop systematic internal planning representations for content or behavior attributes during pre-training, with structure attributes requiring additional reinforcement through fine-tuning.
@@
Figure 4: Layer-wise attribute prediction dynamics. Structural attributes peak mid-layers, content attributes emerge late, and behavioral attributes stabilize early. Layer-wise probing reveals hierarchical organization of planning capabilities, with progressive refinement shaping final outputs.
@@
We investigate pattern saliency by varying the hidden size of two-layer MLP probes and measuring their average performance across model layers. Performance plateaus before hidden size $128$ across all datasets, with larger sizes that can even lead to overfitting, indicating pattern saliency. The results can also indicate saliency differences across attributes: structure attributes (Fig. 3a,  3b) reach converges start at around hidden size $16$, content attributes (Fig. 3c,  3d) plateau around $32$, and behavior attributes (Fig. 3e3f) plateus at around 8, suggesting a hierarchy of representation complexity where behavioral patterns are most readily accessible, structural patterns require moderate complexity to capture, and content patterns demand the most sophisticated probe architectures. The consistent pattern across different model scales and architectures points to fundamental organizational principles in language model representations. This suggests that emergent planning capabilities may be an inherent property of large language models rather than an artifact of specific architectures or training procedures.
@@
We conduct layer-wise probing analysis (with hidden sizes optimized per layer) to understand how different attributes emerge through model layers. The results reveal distinct accumulation patterns for each attribute type. Structure attributes (Fig. 4(a),  4(b)) show weak performance in early layers, peak in middle layers, and partially diminish in final layers, suggesting a gradual accumulation followed by refinement. Content attributes (Fig. 4(c),  4(d)) peak in later layers, either through sudden late-layer emergence or gradual accumulation, indicating their reliance on higher-level semantic processing. Behavior attributes (Fig. 4(e),  4(f)) demonstrate uniform distribution across layers except for the initial few, suggesting they are fundamental properties encoded early in the model. These layer-wise patterns reveal that (1) different aspects of planning emerge through distinct computational paths, (2) the hierarchical nature of planning, from basic behavioral patterns to complex structural decisions, is reflected in the layer-wise organization, and (3) the emergence of these patterns through progressive transformations, rather than from initial embeddings alone, indicates that planning capabilities arise from learned computational processes rather than simple statistical correlations.
@@
(a) Character choice prediction (TinyStories).
(b) Answer confidence prediction (MedMCQA).
Figure 6: U-shaped planning dynamics during generation. Probing at equidistant positions (character choice, answer confidence) shows three-phase patterns: high accuracy in early segments (global planning intent), mid-segment decline (local token focus), and late-stage recovery (contextualized refinement). This suggests models first outline global attributes, then refine locally, before finalizing coherent plans.
Figure 7: Implicit-explicit planning discrepancy. Models struggle to explicitly predict their own response lengths, with base models showing near-zero Spearman correlation and most fine-tuned models achieving only marginal gains. This gap implies limited introspective awareness despite underlying capability.
@@ 5.1 Planning Ability Scales with Model Size
We analyze how emergent response planning scales across different model sizes using four model families: LLama-2-chat (7B, 13B, 70B), Llama-3-Instruct (8B, 70B), Qwen-2-Instruct (7B, 72B), and Qwen-2.5-Instruct (1.5B, 32B, 72B). Using grid search over layers and hidden sizes, we identify optimal configurations and evaluate models on UltraChat and TinyStories datasets, focusing on structure and content attributes. We exclude base models as the relatively small models have short context which limit few-shot prompts, while the same prompts fail to effectively prompt larger base models to follow instructions. We omit the behavior attribute type as larger models tend to give correct answers consistently, making it difficult to obtain balanced data for analysis.
Results shown in Fig. 5 exhibit two key insights: (1) within each model family, larger models demonstrate stronger planning capabilities, and (2) this scaling pattern does not generalize across different model families, suggesting that other factors like architectural differences also influence planning behavior.
@@ 5.2 Evolution of Planning Representations During Response Generation
We analyze how planning features evolve during generation by probing at different positions in the response sequence. For each response, we collect activations from the first token up to the token before attribute-revealing keywords (e.g., animal words in story character selection tasks) or throughout the entire sequence for tasks requiring external ground-truth labels (e.g., answer confidence tasks). We divide these positions into equal segments and apply probes previously trained with in-dataset settings at each division point. We conduct experiments on two datasets: TinyStories for character choice prediction and MedMCQA for answer confidence prediction. Results in Fig. 6 reveal a distinctive pattern: probing accuracy is high initially, decreases in the middle segments, and rises again toward the end. This pattern suggests a three-phase planning process: (1) initial phase with strong planning that provides an overview of the intended response; (2) middle phase with weaker planning, characterized by more local, token-by-token generation; (3) final phase with increased planning clarity as accumulated context makes the target attributes more apparent.
5.3 Gap Between Probing and LLMs’ Self-Predicted Results
We investigate how the models’ ability to predict their own response attributes compare to probe-based predictions with the UltraChat dataset on response length task. For fine-tuned models, we prompt: ”Estimate your answer length in tokens using [TOKENS]number[/TOKENS], then provide your answer.” For base models, we provide few-shot examples with pre-calculated lengths. To evaluate self-prediction accuracy, we compare the predicted token count collected in a separated run against the length of the previous-collected model’s greedy-decoded response.
As shown in Fig. 7, base models achieve near-zero Spearman correlation when predicting token lengths, even with examples. While fine-tuned models perform marginally better, there remains a substantial gap between models’ direct predictions and probe-based predictions. This gap suggests that models encode more planning information in their hidden representations than they can explicitly access during token-by-token generation, indicating a discrepancy between implicit planning capabilities and explicit self-awareness.
@@ 6 Discussion
6.1 Measuring Emergent Response Planning Under Sampling
In this study, we focus on the greedy-decoded responses of LLMs, using greedy decoding to derive deterministic probe labels $\hat{g}_{i}=g(\mathbf{y}_{i})$ for representations $\mathcal{H}_{i}=\{\mathbf{H}^{l}_{\mathbf{x}_{i}}\}^{L}_{l=1}$. But when generalizing to sampling settings, while greedy decoding simplifies sampling approximation by reflecting the LLM’s most probable output, this approach may not fully capture sampling nuances. We propose two potential ways for improvement:
@@ 6.1 Measuring Emergent Response Planning Under Sampling
Our findings on LLMs’ emergent response planning suggest several practical applications. First, anticipating response length and complexity could enable dynamic resource allocation in LLM systems. Second, detecting low confidence or inaccuracies early during generation might trigger interventions like early stopping or retrieval-augmented refinement. Third, predictive capabilities could improve user interaction: predicting reasoning complexity might facilitate task decomposition in multi-step problems, while forecasting response characteristics could enable precise progress estimation during interactive tasks. These possibilities underscore the need to develop robust probing methods for deployed LLMs.

### v2

@@ Abstract
In this work, we argue that large language models (LLMs), though trained to predict only the next token, exhibit emergent planning behaviors: their hidden representations encode future outputs beyond the next token. Through simple probing, we demonstrate that LLM prompt representations encode global attributes of their entire responses, including structure attributes (e.g., response length, reasoning steps), content attributes (e.g., character choices in storywriting, multiple-choice answers at the end of response), and behavior attributes (e.g., answer confidence, factual consistency). In addition to identifying response planning, we explore how it scales with model size across tasks and how it evolves during generation. The findings that LLMs plan ahead for the future in their hidden representations suggest potential applications for improving transparency and generation control.
@@ 1 Introduction
Figure 1: Illustration of probing LLMs for emergent response planning. After processing a prompt, hidden representations ($\mathbf{H}$) are extracted from the LLM’s layers. MLP probes analyze these representations to predict future global attributes, including structure (length), content (“Fox”), and behavior attributes (answer confidence). LLMs can anticipate their future outputs—like the “Fox” appearing at $t_{a}$ or the final length at $t_{N}$—long before generation.
In this work, we argue that LLMs, though trained to predict only the next token, display emergent planning behaviors: their hidden representations encode their future outputs beyond just the next token. Specifically, we observe that LLM prompt representations encode interesting global attributes of their upcoming responses. We call this phenomenon response planning and classify these global attributes into three categories: structure attributes (e.g., response length, reasoning steps), content attributes (e.g., character choices in storywriting, multiple-choice answers at the end of response), and behavior attributes (e.g., answer confidence, factual consistency).
We empirically identify response planning by training simple probes on LLM prompt representations to predict the global attributes of their upcoming responses. We find that these probes achieve non-trivial prediction accuracy, providing strong evidence that LLMs plan at least part of their entire response in advance when they read the prompt. Through further ablations, we find that planning abilities positively scale with model size, peak at the beginning and end of responses, share certain planning patterns across models, and exceed their self-verbalized awareness.
The contribution of our work is two-fold: (1) To our best knowledge, we introduce the first formal definition and framework of emergent response planning in LLMs. (2) We demonstrate empirically that LLMs perform emergent response planning through systematic probing experiments across various attributes types and tasks, and investigate their properties. These findings shed light on LLMs’ internal mechanisms and suggest novel approaches for predicting and controlling outputs pre-generation, potentially enhancing model controllability.
@@ 2 Related Work
Prior works exploring response planning in LLM. Previous studies have examined whether LLMs can anticipate beyond the next token. Future Lens (Pal et al., 2023) models token distributions beyond the immediate next token using linear approximation.  (Geva et al., 2023) studies how LLMs retrieve factual associations during generation, while  (Men et al., 2024) extends this to Blocksworld planning, suggesting LLMs consider multiple planning steps simultaneously.  (Pochinkov et al., 2024) finds that tokens at context-shifting positions may encode information about the next paragraph.  (Wu et al., 2024) hypothesizes LLMs’ lookahead capability and tests two mechanisms—pre-caching and breadcrumbs—in a myopic training setting.
While prior works examine relatively narrow aspects like predictions several tokens ahead or knowledge retrieval in specialized scenarios, our work delves deeper to reveal the broader response planning landscape of LLMs. We provide the first formal definition of response planning in LLMs, investigate comprehensive planning attributes, and demonstrate planning capabilities across diverse real-world tasks.
@@ 3.1 Probing for Future Responses
a Response length prediction.
c Character choices prediction.
e Answer confidence prediction.
b Reasoning steps prediction.
d Multiple-choice answers prediction.
f Factual consistency prediction.
g Example fitting results for response length prediction.
h Example fitting results for reasoning steps prediction.
@@ 3.2 Probing Setup
To study response planning in LLMs, we first design tasks $T=(p(\mathbf{x}),g(\mathbf{y}))$, consisting of a prompt distribution $p(\mathbf{x})$ eliciting key response attributes of interest $g(\mathbf{y})$ as probing targets. Next, we introduce the data collection pipeline for training probes.
Task design. The studied response attributes must be global, meaning they cannot be determined from the first response token and should ideally be distributed across the entire response. We focus on six tasks that elicit response attributes across three categories: structure, content, and behavior.
@@
d Multiple-choice answers.
@@ 3.3 Experimental Details
We experiment with both instruction-tuned models (Llama-2-7B-Chat, Llama-3-8B-Instruct, Mistral-7B-Instruct, and Qwen2-7B-Instruct) and their corresponding base models (Llama-2-7B, Llama-3-8B, Mistral-7B, and Qwen2-7B). See Appendix A.1 for model details.
@@ 4 Experimental Results
Insight 1: Models present emergent planning on structure, content, and behavior attributes, which can be probed with high accuracy (Fig. 2).
Our in-dataset probing experiments (where probes are trained and tested on different splits of the same prompt dataset) reveal that both base and fine-tuned models encode structure, content, and behavior attributes, with fine-tuned models showing superior performance. For structural attributes (response length and reasoning steps; Fig. 2a,  2b), fine-tuned models exhibit strong linear correlations with ground truth, clustering around $y=x$ (with example fitting results shown in Fig. 2g,  2h), while base models show weaker but positive correlations. For content and behavior attributes (character choices, multiple-choice answers, answer confidence, and factual consistency; Fig. 2c,  2d,  2e,  2f), both model types demonstrate robust classification performance above random baselines. These findings also suggest that models develop systematic internal planning representations for content and behavior attributes during pre-training, with structure attributes requiring additional reinforcement through fine-tuning.
@@
Figure 5: Layer-wise attribute prediction dynamics. Six subplots (one per task): Y-axis shows eight models; X-axis traces layer-wise progression (early → late). Heatmap colors indicate absolute performance (0–1, lighter = higher); black curves show row-normalized relative capability trends. Key dynamics: Structure attributes peak mid-layers, content attributes exhibit varied emergence timelines but consolidate in later layers, and behavior attributes stabilize early. Layer-wise probing reveals hierarchical organization of planning capabilities, with progressive refinement shaping final outputs.
@@
We investigate pattern saliency by varying the hidden size of two-layer MLP probes and measuring their average performance across model layers. Performance plateaus before hidden size $128$ across all datasets, with larger sizes that can even lead to overfitting, indicating pattern saliency. The results can also indicate saliency differences across attributes: structure attributes (Fig. 4a,  4b) converges around hidden size $16$, content attributes (Fig. 4c,  4d) plateau around $32$, and behavior attributes (Fig. 4e4f) plateaus at around 8, suggesting a hierarchy of representation complexity where behavioral patterns are most readily accessible, structural patterns require moderate complexity to capture, and content patterns demand the most sophisticated probe architectures. The consistent pattern across different model scales and architectures illustrates fundamental organizational principles in language model representations, suggesting that emergent planning is an inherent property of large language models rather than an artifact of specific architectures or training procedures.
@@
We conduct layer-wise probing analysis (with hidden sizes optimized per layer) to understand how different attributes emerge through model layers. The results reveal distinct accumulation patterns for each attribute type. Structure attributes (Fig. 5(a),  5(b)) show weak performance in early layers, peak in middle layers, and partially diminish in final layers, suggesting a gradual accumulation followed by refinement. Content attributes (Fig. 5(c),  5(d)) peak in later layers, either through sudden late-layer emergence or gradual accumulation, indicating their reliance on higher-level semantic processing. Behavior attributes (Fig. 5(e),  5(f)) demonstrate uniform distribution across layers except for the initial few, suggesting they are fundamental properties encoded early in the model. These layer-wise patterns reveal that (1) different aspects of planning emerge through distinct computational paths, (2) the hierarchical nature of planning, from basic behavioral patterns to complex structural decisions, is reflected in the layer-wise organization, and (3) the emergence of these patterns through progressive transformations, rather than from initial embeddings alone, indicates that planning capabilities arise from learned computational processes rather than simple statistical correlations.
@@
(a) Answer confidence prediction (MedMCQA).
(b) Character choice prediction (TinyStories).
Figure 7: Planning dynamics during generation. Probing at equidistant positions (answer confidence, character choice) shows three-phase patterns: high accuracy in early segments (global planning intent), mid-segment decline (local token focus), and late-stage recovery (contextualized refinement). This suggests models first outline global attributes, then refine locally, before finalizing coherent plans.
@@ 5.1 Planning Ability Scales with Model Size
We analyze how emergent response planning scales across different model sizes using four model families: Llama-2-chat (7B, 13B, 70B), Llama-3-Instruct (8B, 70B), Qwen-2-Instruct (7B, 72B), and Qwen-2.5-Instruct (1.5B, 32B, 72B). Using grid search over layers and hidden sizes, we identify optimal configurations and evaluate models on UltraChat and TinyStories datasets, focusing on structure and content attributes. We exclude base models as smaller models have short context which limit few-shot prompts, while the same prompts fail to effectively prompt larger base models to follow instructions. We omit the behavior attribute type as larger models tend to give correct answers consistently, making it difficult to obtain balanced data for analysis.
Fig. 6 exhibit two key insights: (1) within each model family, larger models demonstrate stronger planning capabilities, and (2) this scaling pattern does not generalize across different model families, suggesting that other factors like architectural differences also influence planning behavior.
(a) Response length prediction.
(b) Character choices prediction.
Figure 8: Gap between probed and verbalized results. Both tasks reveal a systematic gap between verbalized self-estimates (Self-Estimate) and probe-based predictions (Probing). Base models exhibit near or worse-than-random accuracy in self-estimation, while fine-tuned models achieve only marginal gains. This gap demonstrates that models encode richer planning information in hidden representations than they can explicitly access during generation, revealing a divide between implicit planning and explicit self-awareness.
@@ 5.2 Evolution of Planning Representations During Response Generation
We analyze how planning features evolve during generation by probing at different positions in the response sequence. For each response, we collect activations from the first token up to the token before attribute-revealing keywords (e.g., animal words in story character selection tasks) or throughout the entire sequence for tasks requiring external ground-truth labels (e.g., answer confidence tasks). We divide these positions into equal segments and apply probes previously trained with in-dataset settings at each division point. We conduct experiments on two datasets: TinyStories for character choice prediction and MedMCQA for answer confidence prediction. Fig. 7 reveal a distinctive pattern: probing accuracy is high initially, decreases in the middle segments, and rises again toward the end. This pattern suggests a three-phase planning process: (1) initial phase with strong planning that provides an overview of the intended response; (2) middle phase with weaker planning, characterized by more local, token-by-token generation; (3) final phase with increased planning clarity as accumulated context makes the target attributes more apparent.
5.3 Gap Between Probed and Verbalized Results
We investigate whether LLMs can self-estimate their response attributes when explicitly prompted, and compare these verbalized results—self-predictions obtained via direct prompting—against probe-based predictions. Experiments focus on two tasks: response length prediction (Ultrachat) and character choice (TinyStories). For verbalized predictions, we prompt LLMs in separate runs to self-estimate attributes (e.g., “Estimate your answer length in tokens using [TOKENS]number[/TOKENS]” for tuned models, or via few-shot examples with pre-calculated lengths for base models). Self-estimation accuracy is evaluated by comparing these outputs against actual greedy-decoded response attributes.
Fig. 8 reveals a systematic gap: base models exhibit near or worse-than-random accuracy in self-estimation, while fine-tuned models improve only marginally, remaining far inferior to probe-based methods. This suggests models encode more planning information in hidden representations than they can explicitly access during generation, highlighting a divide between implicit planning and explicit self-awareness.
@@ 6 Discussion
6.1 Emergent Response Planning under Sampling
In this study, we consistently use greedy decoding to derive deterministic probing labels $\hat{g}_{i}=g(\mathbf{y}_{i})$ for representations $\mathcal{H}_{i}=\{\mathbf{H}^{l}_{\mathbf{x}_{i}}\}^{L}_{l=1}$. But when generalizing to sampling settings, while greedy decoding simplifies sampling approximation by reflecting the LLM’s most probable output, this approach may not fully capture sampling nuances. We propose two potential ways for improvement:
@@ 6.1 Emergent Response Planning under Sampling
6.2 Defining Planning and Addressing Spurious Correlations
A crucial consideration in defining and measuring “planning” is the potential for spurious correlations, particularly “first-token shortcuts.” An example illustrates this: if a model is prompted in either French or English to provide a ’yes/no’ answer in the same language, a probe analyzing the first token’s activations might predict the final answer (“oui/non” vs. “yes/no”) simply by detecting the language of the initial token. While this shows a correlation, it doesn’t necessarily prove long-range planning in the sense of anticipating specific future content beyond what’s implied by immediate context like language choice.
Our study addresses this by (1) defining planning as the encoding of long-term attributes independently of the immediate next token, and (2) designing prompts to actively block these shortcuts. We define planning such that the hidden representations at the first token should encode both next-token information and long-term attributes, ensuring these two information types are independent – the long-term attribute shouldn’t be directly inferable from the very next token. To implement this in our experiments, we used prompt engineering. For instance, in multiple-choice tasks, we instructed models to first provide an analysis before stating their final answer. This ensures that the initial tokens (the analysis) do not inherently reveal the target attribute (the final choice), thereby helping to isolate genuine planning signals from simpler, shortcut correlations. This methodological approach is vital for ensuring that probes detect true emergent planning rather than just correlated input features.
@@ 6.2 Defining Planning and Addressing Spurious Correlations
Our findings on LLMs’ emergent response planning suggest several practical applications: (1) Pre-generation resource allocation optimization: Probing pre-generation representations allows proactive allocation of computational resources based on anticipated response complexity and length, enhancing inference efficiency during dynamic workloads. (2) Early-error detection: Early detection of behavioral attributes like low confidence could enable corrective interventions (e.g., retrieval-augmented refinement) before errors propagate. Predictive awareness of content attributes (e.g., key entities or argument trajectories) might enable real-time compliance checks with safety guidelines or domain constraints. (3) Novel user interaction paradigms: Predicting reasoning complexity could guide task decomposition for multi-step problems, while predicting response characteristics could improve progress indicators in interactive settings. These possibilities highlight the need for robust probing methods in deployed LLMs.